Subject: APO 3.5m users committee minutes, June 2014
From: Suzanne Hawley
Submitted: Fri, 11 Jul 2014 23:13:06 -0700 (PDT)
Message number: 1337
(previous: 1336,
next: 1338
up: Index)
APO 3.5-m Users Committee Phonecon, 6/23/14
Attending: Suzanne Hawley, Mark Klaene, John Wilson, Remy Indebetouw, Joe
Huehnerhoff, Scott Anderson, Rachel Kuzio de Naray, Bill Ketzeback, John
Bally, Jon Holtzman, Michael Strauss, Joanne Hughes Clark, and Bruce
Gillespie
Note on participants - Suzanne announced that she has invited John Wilson
(UVa) to attend the Users Committee meetings as he is now the Instrument
Scientist for the 3.5-m telescope.
**********************************
User feedback and comments from institutional representatives:
o Oklahoma - No report.
o Colorado - John Bally had nothing new to report.
o Georgia State - Rachel Kuzio de Naray had nothing new to report.
o Princeton - Michael Strauss said everything has been fine. He will be
soon sending Suzanne some internal Princeton time trades for the quarter.
o Seattle Univ. - Joanne Hughes Clark reported that after climbing the
learning curve with ARCSAT, everything has been fine.
o NMSU - Jon Holtzman had nothing new to report.
o Washington - Scott Anderson had nothing new to report.
**********************************
Discussion of telescope/instruments report:
Mark highlighted the recent utility power problems at the site. During a
power outage, when commercial power was restored it was not "clean" and
damaged a transfer switch at the site that is involved in the automated
system to use the backup generator. This is a major system component that
will be replaced soon, possibly during the 3.5m shutdown. In the
meantime, we will need to switch from commercial to generator power
manually if the need arises, with some attendant risks. The rest of his
comments were largely from statements contained in the written report,
which follows:
*************************
3.5-m Telescope and Instruments Highlights
05/16/14 through 06/17/14
0) Overview
Generally another productive period with variable cloudiness, little
precipitation and with typical spring wind and dust events in late May.
Current forest fire danger is still "Very High". Three major site
infrastructure problems arose this period; the new computer room HVAC unit
failed and took out 3 components, an underground water leak was
discovered, and a surge on the incoming utility power took out the
automatic transfer switch. All but the automatic transfer switch have
been repaired. The part is on order and expected in the upcoming weeks.
We are currently vulnerable to a site power outage if utility power drops
for more than 15 minutes with no one on site. Once the part comes in
power will be out on site for about a day and may cause disruption to part
of the night schedule.
A number of training sessions were held this period. Indiana, Washington,
Middlebury and JHU all sent people to the site for general 3.5m
certification and a GIFS training session was also held on site by the
GSFC folks.
1) Telescope
Telescope has been working well except for occasional az oscillations
during or at the end of slews. The area seems to be in 1 part of the sky
and we are trying to find the compromise between these slew issues and
tracking performance. We are also looking at mechanical issues as well as
issues that differ between the old and new amplifiers. The new TCC
development is also progressing nicely with continued testing on the
telescope. Numerous pre-shutdown summer maintenance items have been
accomplished.
2) Instruments
DIS has been operating normally.
Agile is operational with no significant problems.
TripleSpec had no operational problems this period.
SPICam variable dark current has not re-appeared after ion pump work last
period.
Echelle has been operational.
NIC-FPS continues to operate in shared-risk observing mode due to
occasional image corruption issue. NFS less than 9 is operational.
GIFS is operational. The small LN2 hold time designed in by GFSC
continues to be a problem. The training session was attended by several
staff and observers and was successful with only a couple documentation
and software issues.
3) ARCTIC
The instrument development is proceeding on schedule. Once the final
report is published and reviewed a go-ahead is expected on the final
optics design which has received tentative approval.
4) Summer shutdown
There are 2 shutdowns scheduled this summer. The first is from July
7th to July 24th for M2 and M3 mirror re-coating, mechanical maintenance,
electronic system upgrades and the majority of instrument servicing. A
second shutdown will be from August 8th to August 17th for the new TCC
upgrade. We expect the telescope to be operational for much of the time
during the second shutdown as this is necessary to test the new TCC.
5) Guider matching scripts have been rewritten due to time reporting
change in headers, and are now available on newton
**********************************
GIFS training and future use - Ketzeback:
Bill reported that some users came to APO for GIFS training along with APO
staff. The training was led by Carol Grady from GSFC. It generally went
well and encompassed almost all aspects of GIFS operations. John Bally
noted that the instrument has good capabilities, although the FOV is
small. For future usage of GIFS, Suzanne noted that at present only a
couple of institutions have users that want to schedule GIFS on the
telescope. Since the instrument has a relatively short LN2 hold time, we
plan to keep the instrument warm until its next scheduled run in
September. If anyone needs it sooner, we need to be given adequate
notice. Suzanne added that the plan will likely be to block-schedule GIFS
in future quarters unless the demand for it increases.
**********************************
ARCSAT observing and feedback - Ketzeback:
Bill said that we are about 6 weeks into the shared-risk observing period
with ARCSAT. Operations are generally going well, although there is a
substantial learning curve for the not-like-TUI user interface. Joanne
said that she used ARCSAT for about a week with her students, and the only
issue was getting the survey camera to cool to operational temperatures
quickly. Bill added that he is hoping that as users from our partner
institutions become proficient with ARCSAT usage, they can serve as the
local gurus to help train other users. Suzanne mentioned that she has
clarified a policy: PhD graduate students can propose to use ARCSAT on
their own, but all other students (undergrads, masters, etc.) must be
accompanied by a PhD astronomer for (remote) training. The shared-risk
observing with ARCSAT is scheduled through September. There will be a
call for proposals in August for observing programs in quarter 4.
On technical topics, there is now a way to bypass the 3.5-m interlock so
that users can run ARCSAT when the 3.5-m is closed (this is a
weather-safety interlock). Suzanne mentioned that we ordered the
Johnson-Cousins filter set, and we found the Stromgren set. We also have
several H-alpha filters. Users should contact Bill Ketzeback if they need
special filters for an ARCSAT run. Jon asked if we have u ser wiki page
for ARCSAT. Mark said that we need to start one, and Jon offered to put
the information on his APO wiki page. In addition, users should get on
the ARCSAT mailers, which consist of 05m-obs (where the night logs go) and
05m-gen (the general mailer for ARCSAT). You can sign up for these
mailers by going to the main APO web page and look for the link to the
ARCSAT pages. Lastly, Suzanne gave congratulations to Bill and Joe for
all the effort they have put into making ARCSAT a user-friendly and
capable telescope.
**********************************
June 16 Board meeting recap - Hawley:
Suzanne and Rene held a telephone meeting with the ARC Board of Governors
to discuss the options for future leases of telescope time. The situation
has developed that we are anticipating being oversubscribed between the
leasing requests and ARC partner allocations. The proposal is to ask ARC
partner institutions to forgo taking the full measure of their telescope
time in the near term, thus allowing leasing partners to help continue to
build up our financial contingency. The Board agreed with this approach,
and we will therefore be signing some new leases this summer. The Board
also decided to raise the leasing cost, as a response to the strong demand
for 3.5m telescope time.
**********************************
Preliminary ideas about a new spectrograph - Hawley
Suzanne is soliciting ideas for a new optical spectrograph for the 3.5-m,
to replace DIS in the next few years. The first step is to decide what
features we would want with a new instrument (e.g., IFU, time-resolved,
DIS-like attributes, etc.). John Wilson said that he would help organize
the push for the definition of a new optical spectrograph. There was also
some discussion on rekindling the project to get a fiber run from the
3.5-m to the APOGEE H-band spectrograph.
**********************************
Miscellaneous - Hawley
The apo35-general mailing list is going to be migrated from its Princeton
server, where it has been managed by Michael Strauss for nearly 20 years,
to an APO-hosted mailing list. Stay tuned over the summer for more
information.
ACTION ITEMS:
[open from previous meeting]:
none
[new from this meeting]:
none
**********************************
Next meeting: The next Users Committee phonecon will be on Monday 25
August (**not** 18 August as announced last month) at 8:30 AM Pacific
Time. We will skip July due to the 3.5m shutdown. The agenda and other
materials will be sent to the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1337 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO